# DMin: Scalable Training Data Influence Estimation for Diffusion Models

### v3

@@ 4 Experiments
Datasets. For the conditional diffusion model, we combine six datasets from Huggingface and randomly select 80% of the data samples as the training dataset, resulting in 9,288 pairs of images and prompts. Due to page limitations, we list three datasets used for evaluation in Table 1: (1) Flowers, which includes 162 training pairs of flower images and corresponding descriptive prompts in our experiments, accounting for only 1.74% of the training dataset. (2) Lego Sets: This subset consists of 40 training pairs, where each image represents a Lego box accompanied by a description of the box, accounting for only 0.43% of the training dataset. (3) Magic Cards, which contains magic card images from scryfall with captions generated by Fuyu-8B [3] and BLIP [20]. For unconditional diffusion models, we mainly focus on MNIST and CIFAR-10. We include a detailed explanation of datasets [^5] MohamedRashad/midjourney-detailed-prompts [^6] HighCWu/diffusiondb_2m_first_5k_canny777svjack/pokemon-blip-captions-en-ja in Appendix 7.2.
Models. For conditional text-to-image diffusion model, we use three different models: (1) SD 1.4 with LoRA [^1] CompVis/stable-diffusion-v1-4, (2) SD 3 Medium with LoRA333stabilityai/stable-diffusion-3-medium-diffusers and (3) SD 3 Medium33footnotemark: 3 (Full parameters). For unconditional diffusion model, we conduct experiments on two Denoising Diffusion Probabilistic Models (DDPM) trained on MNIST2221aurent/ddpm-mnist and CIFAR-10444google/ddpm-cifar10-32. The detailed settings of models are included in Appendix 7.1. We fine-tune models on the combined training dataset mentioned above and evaluate them on the testing dataset. During gradient collection, we collect only the gradients of the parameters in the LoRA components for the LoRA-tuned model, whereas for the fully fine-tuned model, we collect the gradients of all parameters (Figure 2(b)).
@@
Table 2: Average detection rates of top-k most influential training data samples. Detection rate=# Samples from Same Subset among Top-k Training $\text{Detection rate}=\frac{\text{\# Samples from Same Subset among Top-k Training Samples}}{k}$ where $k=\{5,10,50,100\}$, indicating the average proportion of samples from the same subset appearing in the top-k influential samples. “Ours (w/o Comp.)” indicates that the gradient vectors are not compressed, while “w/o Norm.” signifies that the gradient vectors are not normalized. “Excatly” denotes exact inner product computation. The results for LiSSA, DataInf and D-TRAK on SD3 Medium (Full) are omitted due to the requirement of over hundreds of TB of cache. Moreover, at this scale, it is impractical for LiSSA and DataInf to approximate the Hessian inversion and for D-TRAK to compute an extremely large random projection matrix.
Baselines We compare the proposed DMin against the following baselines: (1) Random Selection: Assigns an influence score to each training sample randomly. (2) SSIM [4] Structural Similarity Index Measure (SSIM) between the training image and the generated image. (3) CLIP Similarity [30]: Cosine similarity of embeddings computed by CLIP between the training image and the generated image. (4) LiSSA [1]: A second-order influence estimation method that uses an iterative approach to compute the inverse Hessian-vector product. (5) DataInf [18]: An influence estimation method based on a closed-form expression for computational efficiency. We also evaluate a variant of DataInf where the Hessian Inversion matrix is replaced with an identity matrix. (6) D-TRAK [27]: A first-order influence estimation method extended from TRAK [28]. (7) Journey-TRAK [11]: An estimation method focusing on the sampling path in diffusion models. For the proposed DMin, we evaluate DMin under different scenarios, including exact estimation of influence scores for each training sample and KNN-based approximate searches for the top-$k$ most influential samples. Additionally, we experiment with varying compression levels: no compression, and $v=\{2^{12},2^{16},2^{20}\}$. The detailed information of baselines are reported in Appendix 8.1.
KNN. We utilized hierarchical navigable small world (HNSW) algorithm [25] for KNN in our experiment, and provide the results of ablation study in Appendix 8.
@@ 4.1 Performance on Conditional Diffusion Models
We begin by generating images using the prompts from the test set of each subset – Flowers, Lego Sets, and Magic Cards. For each test prompt and its generated image, we estimate the influence score for every training data sample and select the top-$k$ most influential training samples with the highest influence score. We then calculate the detection rate as Detection Rate=# Samples from Same Subset among top-k Training $\text{Detection Rate}=\frac{\text{\# Samples from Same Subset among top-$k$ Training Samples}}{k}$.
Results. We report the average detection rate for each test set of subsets in Table 2. Compared to the baselines, our proposed DMin obtains the best performance on all subsets. Compared to the baselines, our proposed DMin achieves the best performance across all subsets. The detection rates for top-50 and top-100 on Lego Sets are lower because the Lego Sets training dataset contains only 40 samples (0.43% of the total). Across all subsets and different values of $k$, $v=2^{16}$ achieves the best performance in most cases, whether using KNN or exact inner product computation. Additionally, compared to our method without compression, removing normalization substantially decreases performance, confirming that normalization mitigates the instability of gradients in extremely deep models. Interestingly, KNN search often outperforms exact inner product computation in our experiments across all models and subsets. This improvement is likely due to KNN’s ability to approximate the search process, capturing a broader and more representative subset of neighbors.
@@
Table 3: Storage requirements for caching gradients of per-sample and entire dataset (9,288 samples), comparing methods with and without compression across different models. The table reports the storage size and compression Ratios achieved by our method at various compression levels relative to uncompressed gradient storage, with LiSSA and DataInf caching entire gradients without compression.
@@
Table 4: Time cost comparison and speedup relative to our method without compression. The time cost represents the duration required to estimate the influence score for every training sample in the dataset for a single test sample (in seconds per test sample).
@@ 4.2 Time and Memory Cost
Time. Table 4 demonstrates the time consumption on estimate the influence score for every training sample in the training dataset for a single test sample. The gradient computation and caching times for two LoRA-tuned models are nearly identical due to the small model size across different methods: (1) SD 1.4 (LoRA): around 8 GPU hours, (2) SD 3 Medium (LoRA): around 24 GPU hours, and (3) SD 3 Medium (full): 330 GPU hours. Additionally, the index construction process only takes a few minutes. Our proposed methods demonstrate substantial efficiency improvements, particularly with KNN search. For instance, on the smallest subset—Lego Sets, which contains only 21 test samples—estimating the influence score for the entire training dataset takes 17 hours with LiSSA, 7 hours with DataInf (Hessian Inversion), and 2 hours with D-TRAK. In contrast, our method with $v=2^{12}$ and $k=5$ requires only 0.084 seconds, and even for $k=100$, it takes only 0.69 seconds while achieving the best performance.
@@ 5 Related Work
Influence estimation has been a critical area of research in understanding the impact of individual training samples on machine learning models [32, 28, 37, 6]. Early work by Koh and Liang [16], Agarwal et al. [1] proposed second-order Hessian-based methods to approximate the effect of a training sample. However, approximating a Hessian inversion becomes computationally prohibitive for large-scale datasets and modern models containing billions of parameters. To address this issue, some studies proposed first order approaches for influence estimation [29, 28]. However, even with first-order methods, scaling to large datasets still encounter storage challenges. For example, storing the gradient of a 2B diffusion model for 10,000 data samples across 10 timesteps requires over 700 TB of storage.
To reduce the storage and computational demands, some studies leverage dimension reduction techniques [28, 27, 11, 14], such as random projection. However, while random projection can substantially reduce the dimension of gradient vector, the projection matrix itself becomes a scalability bottleneck in large models. For instance, in a model with 2B parameters, a projection matrix mapping gradients to a compressed dimension of 32,768 would require over 500 GB of storage. These constraints highlight the need for more efficient and scalable approaches.
@@ 6 Conclusion
In this paper, we introduce DMin, a scalable framework for estimating the influence of training data samples on images generated by diffusion models. The proposed DMin scales effectively to diffusion models with billions of paramters by substantially reducing storage requirements from hundreds of TBs to MBs or KBs for SD 3 Medium with full parameters. Additionally, DMin can retrieve the top-$k$ most influential training sample in 1 second by KNN, demonstrating the scalability of the proposed DMin. Our empirical results further confirm DMin’s effectiveness and efficiency.
@@
In this section, we report the detailed setting and environments for our experiments.
@@
Experimental Environments. Our experiments are conducted on four different types of servers: (1) Servers running Red Hat Enterprise Linux 7.8, equipped with Intel(R) Xeon(R) Platinum 8358 processors (2.60GHz) with 32 cores, 64 threads, 4 A100 80G GPUs, and 1TB of memory. (3) Servers running Red Hat Enterprise Linux 7.8, containing Intel(R) Xeon(R) Gold 6226R CPUs @ 2.90GHz with 16 cores, 32 threads, 2 A100 40G GPUs, and 754GB of memory. (4) A server running Ubuntu 20.04.6 LTS, featuring 2 H100 GPUs, dual Intel(R) Xeon(R) Gold 6438N processors (3.60GHz) with 32 cores, 64 threads, and 1.48TB of memory. To ensure a fair comparison, all experiments measuring time cost and memory consumption are conducted on server 1, while other experiments are distributed across the different server types.
@@ 7.2 Datasets
In this section, we introduce the datasets used on our experiments of conditional diffusion models and unconditional diffusion models.
@@
SD 3 Medium (Full) 0.0001 64 150 $512\times 512$ float32
@@ 8 Ablation Study
Figure 5: Additional visualization for unconditional diffusion model on the MNIST dataset.
@@ 9 Supplemental Visualization for Conditional Diffusion Models
Figure 6: Examples of the top-25 most influential training data samples for the generated image (the 1-st column) on SD 3 Medium with LoRA, shown from high to low influence from left to right.

### v4

@@ 4 Experiments
Datasets. For the conditional diffusion model, we combine six datasets from Hugging Face and randomly select 80% of the data samples as the training dataset, resulting in 9,288 pairs of images and prompts. Due to page limitations, we list three datasets used for evaluation in Table 1: (1) Flowers, which includes 162 training pairs of flower images and corresponding descriptive prompts in our experiments, accounting for only 1.74% of the training dataset. (2) Lego Sets: This subset consists of 40 training pairs, where each image represents a Lego box accompanied by a description of the box, accounting for only 0.43% of the training dataset. (3) Magic Cards, which contains magic card images from Scryfall with captions generated by Fuyu-8B [3] and BLIP [20]. For unconditional diffusion models, we mainly focus on MNIST and CIFAR-10. We include a detailed explanation of datasets55footnotemark: 566footnotemark: 677footnotemark: 7 in Appendix 7.2.
Models. For conditional text-to-image diffusion models, we use three different models: (1) SD 1.4 with LoRA, (2) SD 3 Medium with LoRA and (3) SD 3 Medium (Full parameters). For unconditional diffusion models, we conduct experiments on two Denoising Diffusion Probabilistic Models (DDPM) trained on MNIST and CIFAR-10. The detailed settings of models are included in Appendix 7.1. We fine-tune models on the combined training dataset mentioned above and evaluate them on the testing dataset. During gradient collection, we collect only the gradients of the parameters in the LoRA components for the LoRA-tuned model, whereas for the fully fine-tuned model, we collect the gradients of all parameters (Figure 2(b)).
[…]
@@
Baselines. We compare the proposed DMin against the following baselines: (1) Random Selection: Assigns an influence score to each training sample randomly. (2) SSIM [4]: Structural Similarity Index Measure (SSIM) between the training image and the generated image. (3) CLIP Similarity [35]: Cosine similarity of embeddings computed by CLIP between the training image and the generated image. (4) LiSSA [1]: A second-order influence estimation method that uses an iterative approach to compute the inverse Hessian-vector product. (5) DataInf [18]: An influence estimation method based on a closed-form expression for computational efficiency. We also evaluate a variant of DataInf where the Hessian Inversion matrix is replaced with an identity matrix. (6) D-TRAK [30]: A first-order influence estimation method extended from TRAK [32]. (7) Journey-TRAK [11]: An estimation method focusing on the sampling path in diffusion models. For the proposed DMin, we evaluate DMin under different scenarios, including exact estimation of influence scores for each training sample and KNN-based approximate searches for the top-$k$ most influential samples. Additionally, we experiment with varying compression levels: no compression, and $v=\{2^{12},2^{16},2^{20}\}$. Details of the baselines are reported in Appendix 9.1.
KNN. We use the hierarchical navigable small world (HNSW) algorithm [27] for KNN in our experiments, and we provide the results of an ablation study in Appendix 8.
@@ 4.1 Performance on Conditional Diffusion Models
[…]
Results. We report the average detection rate for each test set of subsets in Table 2. Compared to the baselines, our proposed DMin achieves the best performance across all subsets. The detection rates for top-50 and top-100 on Lego Sets are lower because the Lego Sets training dataset contains only 40 samples (0.43% of the total). Across all subsets and different values of $k$, $v=2^{16}$ achieves the best performance in most cases, whether using KNN or exact inner product computation. Additionally, compared to our method without compression, removing normalization substantially decreases performance, confirming that normalization mitigates the instability of gradients in extremely deep models. Interestingly, KNN search often outperforms exact inner product computation in our experiments across all models and subsets. This improvement is likely due to KNN’s ability to approximate the search process, capturing a broader and more representative subset of neighbors.
Table 3: Storage requirements for caching per-sample and dataset gradients (9,288 samples), comparing compressed and uncompressed methods across models. The table shows storage and compression ratios of our method across levels, with LiSSA and DataInf storing gradients uncompressed.
@@
Table 4: Time cost comparison and speedup relative to our method without compression. Time cost refers to the time to estimate influence scores for all training samples per test sample (in seconds).
@@ 4.2 Time and Memory Cost
Time. Table 4 reports the time required to estimate the influence score for every training sample in the training dataset for a single test sample. The gradient computation and caching times for the two LoRA-tuned models are similar across methods: (1) SD 1.4 (LoRA): around 8 GPU hours, (2) SD 3 Medium (LoRA): around 24 GPU hours, and (3) SD 3 Medium (full): 330 GPU hours. Additionally, the index construction process only takes a few minutes. Our proposed methods demonstrate substantial efficiency improvements, particularly with KNN search. For instance, on the smallest subset—Lego Sets, which contains only 21 test samples—estimating the influence score for the entire training dataset takes 17 hours with LiSSA, 7 hours with DataInf (Hessian Inversion), and 2 hours with D-TRAK. In contrast, our method with $v=2^{12}$ and $k=5$ requires only 0.084 seconds, and even for $k=100$, it takes only 0.69 seconds while achieving the best performance.
@@ 5 Related Work
Influence estimation has been a critical area of research in understanding the impact of individual training samples on machine learning models [37, 32, 42, 6]. Early work by Koh and Liang [16], Agarwal et al. [1] proposed second-order Hessian-based methods to approximate the effect of a training sample. However, approximating a Hessian inversion becomes computationally prohibitive for large-scale datasets and modern models containing billions of parameters. To address this issue, some studies proposed first-order approaches for influence estimation [34, 32]. However, even with first-order methods, scaling to large datasets still encounters storage challenges. For example, storing the gradient of a 2B diffusion model for 10,000 data samples across 10 timesteps requires over 700 TB of storage.
To reduce the storage and computational demands, some studies leverage dimension reduction techniques [32, 30, 11, 14], such as random projection. However, while random projection can substantially reduce the dimension of the gradient vector, the projection matrix itself becomes a scalability bottleneck in large models. For instance, in a model with 2B parameters, a projection matrix mapping gradients to a compressed dimension of 32,768 would require over 500 GB of storage. These constraints highlight the need for more efficient and scalable approaches.
@@ 6 Conclusion
In this paper, we introduce DMin, a scalable framework for estimating the influence of training data samples on images generated by diffusion models. The proposed DMin scales effectively to diffusion models with billions of parameters by substantially reducing storage requirements from hundreds of TBs to MBs or KBs for SD 3 Medium with full parameters. Additionally, DMin can retrieve the top-$k$ most influential training samples in under one second using KNN, demonstrating the scalability of the proposed DMin. Our empirical results further confirm DMin’s effectiveness and efficiency.
Acknowledgments
This work was supported by the National Science Foundation under Award Nos. 2543795, 2247619, and 2413046. We acknowledge Research Computing [29] at RIT for providing computing resources. We also acknowledge the Tufts University High Performance Compute Cluster [33], which was utilized for the research reported in this paper.
@@ Acknowledgments
In this section, we report the detailed experimental settings and environments.
@@
Experimental Environments. Our experiments are conducted on three types of servers: (1) Servers running Red Hat Enterprise Linux 7.8, equipped with Intel(R) Xeon(R) Platinum 8358 processors (2.60GHz) with 32 cores, 64 threads, 4 A100 80G GPUs, and 1TB of memory. (2) Servers running Red Hat Enterprise Linux 7.8, containing Intel(R) Xeon(R) Gold 6226R CPUs @ 2.90GHz with 16 cores, 32 threads, 2 A100 40G GPUs, and 754GB of memory. (3) A server running Ubuntu 20.04.6 LTS, featuring 2 H100 GPUs, dual Intel(R) Xeon(R) Gold 6438N processors (3.60GHz) with 32 cores, 64 threads, and 1.48TB of memory. To ensure a fair comparison, all experiments measuring time cost and memory consumption are conducted on server type (1), while other experiments are distributed across the different server types.
@@ 7.2 Datasets
In this section, we introduce the datasets used in our experiments on conditional diffusion models and unconditional diffusion models.
@@
SD 3 Medium (Full) 0.0001 64 150 $512\times 512$ - - - float32
@@ 8 Ablation Study
9 Additional Analysis
Training order and data augmentation. Our influence score is defined with respect to the final trained parameters $\theta$, rather than intermediate checkpoints. While training order affects the optimization trajectory, its effect is absorbed into the final parameter state used for influence estimation, especially under large-scale training with extensive shuffling. Data augmentation is treated as part of the training distribution; in practice, we report influence at the image level instead of treating each augmented view as a separate training sample.
Timestep subsampling and gradient correlation. Adjacent diffusion timesteps are often highly correlated. DMin accounts for this by subsampling timesteps for influence estimation instead of computing gradients for all $T$ timesteps, as described in the main method. This design reduces computation and storage while preserving performance.
Effect of removing influential samples. We further study whether removing highly influential samples changes model behavior on MNIST. We first train an unconditional DDPM on the full 60,000-image training set and generate 5,000 samples. Using a stroke-thickness metric, we select 500 generated images with the thickest strokes, run top-1000 influence estimation for each image, and deduplicate the retrieved training samples, yielding 14,562 influential samples. After removing these samples and retraining on the remaining 45,438 images, generated digits shift toward thinner strokes, indicating that removing highly influential samples leads to systematic and interpretable changes in generation behavior (Figure 5).
Figure 5: Removing highly influential MNIST training samples systematically shifts generated digits from thick to thinner strokes.
Figure 6: Additional visualizations for unconditional diffusion models on the MNIST dataset.
@@ 10 Supplemental Visualization for Conditional Diffusion Models
Figure 7: Examples of the top-25 most influential training data samples for the generated image (first column) on SD 3 Medium with LoRA, shown from high to low influence from left to right.